Instrument Log 

Picarro C1 CF Lamont, OK

Date

Technician

Comments

4/6/16

Mentor

Instrument mentor on site and replaced the Nafion box for the Picarro.
*Old - S/N: 001 (Model: NAFION-4CH)

*New - S/N: 003 (Model: NAFION-1CH)

2/3/16

MG

12-7-2015 at 15:00 GMT Pump 3 on instruments went bad which caused the pressure gauge on pump 3 to max out and get stuck. 2-3-2016 Instrument mentor sent out a new pressure gauge. I swapped it out and adjusted pressure accordingly. RTS 2-3-2016 20:30 GMT.

1/27/16

MG

Front and rear fan flow meters on the pump box were reading negative. Mentor and I shut down the Picarro at 17:30 GMT we removed the pump box and switched the orientation of the flow meters. this did not fix the problem. RTS 1806 GMT.

12/31/16

MG

The Central Facility lost power due to an ice storm on Sunday 12-27-2015 around 00:00 GMT Before restarting system I replaced the pump box per mentors request. I then Reset the building UPS and started program RTS 12-31-2015 at 17:35 GMT.

12/8/15

MG

Pump 3 on Picarro went bad. Mentor is shipping a new pump rack, which should arrive in January.

12/2/15

MG

Piccarro was installed by mentor at 2200 GMT on 12/2/2015.

7/1/14

KT

Instrument mentor requested that the Picarro be removed and shipped back for repairs. At 1325 GMT removed the *Picarro and took down to shipping/receiving to be shipped out for repairs. RTS at 1340 GMT. Model# ESP-1000, SN# 159-CPVU-CFA047.

Model# ESP-1000-DAS, SN# 164-CFAD5047, WD# 59050.

2/27/14

KT

Pump #4 was not operational. At 1700 GMT removed pump SN# 1/1041919 and installed Pump SN# 1/952959. RTS at 1800 GMT.

11/21/13

KT

The *Picarro was returned from being repaired and needed to be installed. At 1500 GMT the Picarro Cavity Ringdown Spectrometer was installed. Instrument mentor requests that the system be ran and tested for moisture leaks. SN# 164-CFAD5047, WD# 59050, model ESP-1000-DAS. Model# ESP-1000, SN# 159-CPVU-CFA147.

8/27/2013

PAD/KT

At 1435 GMT, found #1 and #2 sample drain line tubes were blocked. Removed blockage with canned air. At 1506 GMT, connected a spare desiccant canister to line labeled CO2 inlet in bottom box of rack. Desiccant canister inlet is open to room air. Observed water in clear tubing coming from CO2 purge out in bottom box. Tubing is connected to external vacuum pump. Removed line, blew out water and reconnected line. This had to be done a couple of times as more water came into tubing when it was reconnected. At this time, following values are shown for Picarro system. CH4=2.033, CO2=408.05, H2O=0.000, Drier flow=5.49, Drier temp=9.64, Drier pressure= -3.67, Fridge temp= 10.87. RTS at 1535 GMT.

3/12/13

PAD

CO2 level is not shown on graph. No flow is shown on Controller/pressure tab for current sampling position. Error is shown: Replace existing "C:\Picarro|Results|20130311|CFADS47-20130311-0518-Log.dat"? Replace or Cancel. Have not seen this error message before and was unsure what program was asking to replace or what it was going to replace it with. Left voicemail with Mentor. Mentor later called and stated file in question was empty. He selected Replace command and program restarted correctly. RTS at 1653 GMT.

2/26/13

PAD

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm. Power was restored at 1920 GMT. All pumps and Picarro system restarted after power was restored. Program log shows WB, warm body temp alarm. At 1400 on 2/27/12, observed program was still showing WB temp alarm. Rebooted computer and program started correctly. RTS at 1415 GMT.

2/20/13

PAD/KT

While troubleshooting NOAA flasks sampler issue, observed Picarro CO2 concentration would rise above normal level when sampling air from 4m level, pump #2. Mentor suspects pump diaphragm has failed allowing room air to be drawn into sample line and requests pump replaced with spare sent to SGP. Shut down all Picarro pumps at 1623 and removed pump rack. R/R pump #2, installed rack and restarted pumps. Old KNF type UN010, SN 1/1041922. New KNF type UN010, SN 1/1195183. Started pumps at 1641 GMT, monitored all pumps for flow and CO2 concentration. All pumps briefly showed higher CO2 concentration due to sample lines being open to room air during pump replacement but quickly returned to normal CO2 levels, ~406 ppm. RTS at 1658 GMT.

1/3/13

KT

Instrument mentor notified that he was not able to connect to the PGS/PICARRO computer. At 1940 GMT there was no response from the PGS/PICARRO computer. Rebooted computer and called instrument mentor who was able to connect to the computer.

12/18/12

KT

The PGS Picarro/Target cylinder pressure had dropped significantly. During inspection a leak was discovered where the regulator connects to the cylinder. Turned gas off, removed regulator and wrapped teflon tape around end. Connected regulator to tank. Leak had stopped. RTS at 1638 GMT.

12/10/12

KT

Instrument mentor requested that the Picarro/Target cylinder tank be removed and replaced. At 2030 GMT removed and replaced the Picarro/Target cylinder. RTS at 2100 GMT. Old SN# CB08983, New SN# CB08814.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County. Power was restored at 2115 GMT after electrical upgrades were complete.

11/13/12

PAD/KT

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound. Power was restored at approximately 1930 GMT after all transformers were replaced. At 2110 GMT, checked flows on Picarro system. Verified program was operating correctly. RTS at 2130 GMT.

9/5/12

PAD

Mentor requests Picarro pump rack replaced and removed pumps sent back to him to be rebuilt. Shut down pumps at 2000 GMT. R/R pump rack assembly. Powered up pumps and adjusted flow as needed. RTS at 2020 GMT.

6/19/12

PAD

Span 1 and span 2 cylinders are below 350 psi. Mentor sent replacement cylinders to site and requests cylinders swapped. Shut off cylinder valves at 1620 GMT. R/R cylinders, checked for leaks at regulators, and purged lines from cylinders to Picarro. Span 1 OLD CB08883, NEW CB09060. Span 2 OLD CB03802, NEW CB09697. Removed cylinders are stored on site for later verification. RTS at 1650 GMT.

3/28/12

PAD

Mentor requests CO2 cylinders used for IRGA calibrations connected to Picarro instrument to verify concentrations. At 1407 GMT, connected cylinder CC235442, (461.2 ppm CO2), to inlet #2 on Picarro instrument. CO2 cylinder was sampled on #2 inlet until 1705 GMT. At that time cylinder was removed and cylinder CC201793 was connected. CO2 cylinder, (459.6 ppm CO2), was sampled on #2 inlet until 1932 GMT. Cylinder was removed and #2 inlet line was attached to instrument. RTS at 1947 GMT.

3/22/12

CAM

Mentor notified Site Ops of abnormalities in the CO data which suggested possible water in the lines. Checked instrument at 1845 GMT and found about 2 of water in the tub where water traps drain to.   Opened the fridge and observed quite a bit of water in the CO water trap, as well as the others, even though all the drain lines were clear and bubbling.   There was also significant moisture in the CO sample lines between the pump outlet and pressure regulator/gauge, and some visible moisture in the line going to the instrument.  Disconnected each section of tubing one at a time and let the water blow out and also blew out with canned air.  Opened up and drained all the water traps in the fridge. RTS at 1920 GMT. Note there has been over 2 inches of rain here this week and high humidity when not raining. Also the fridge temperature has been running around 7C with the thermostat set to max cool.  Notified mentor of findings.

11/23/11

PAD

Picarro target cylinder show 280 psi and needs to be replaced. R/R target cylinder. Reconnected regulator and checked for leaks. Took removed cylinder to shipping dept to be sent to vendor and be refilled. Old CB08814. New CB08983. RTS at 1706 GMT.

8/23/11

PAD

Mentor requests vacuum pump rack for Picarro system replaced with rebuilt pump rack supplied by mentor. Shut down pumps and R/R pump rack. Powered up pumps. RTS at 1933 GMT. Old 1/1026011, 1/1041921, 1/1041922, 1/952959. New 1/1336619, 1/1332023, 1/1336621, 1/1336620.

5/24/11

PAD

Loggernet is not on monitor computer. Tried to start Loggernet but message shows: Unable to start Loggernet program. Loggernet is already running. Stopped data log and shut down program. Rebooted computer. Loggernet started after computer reboot and now able to connect with datalogger. RTS at 1528 GMT.

6/9/2011

JEM

Found pump #2 4M switch in the "OFF" position. Powered the switch back on @1805 GMT. RTS @1815 GMT.

6/8/2011

JEM

The mentor called and said the Picarro refrigerator was getting upwards of 20C. An inspection of the refrigerator found that the entire upper portion of the refrigerator was solid ice. Unplugged the refrigerator and proceeded to thaw it out using a heat gun. Used towels to soak up and discard the water from the melted ice. After the refrigerator was thawed out, the temperature began to slowly drop. RTS @2100 GMT.

4/25/11

PAD

Picarro computer was removed and sent for repairs. It has been returned and needs to be installed. Installed computer and connected all cables and tubes. Powered up Picarro system and notified mentor. Mentor was able to remotely access computer and monitor system. Computer model ESP-1000, SN 159-CPVU-CFR047. RTS at 1625 GMT.

4/11/11

PAD

Notified by mentor he is unable to remotely access Picarro computer. Last data collected was sometime on 4/8/11. At 1913 GMT, observed power indicator ring light on computer was blinking every 4 seconds. Read TS section in manual and pressed power button with no results. Contacted Picarro technical support and was advised to reseat memory card located in computer tray. Removed computer tray from rack and reseated memory card, installed computer and powered up system. No change to computer status. Removed computer tray and took to electronics lab. Reseated card again and looked for obvious cause of problem. Applied power to computer and harddrive would run a few seconds, cooling fans would start, then system would shut down, and repeat. Troubleshooting with vendor and mentor at this time. RTS at 2100 GMT.

12/29/10

PAD

At 2101 GMT, observed errors shown in status log and also spectrometer display: Scan Timeout, Timeout: Measurement. CO2 and Methane graphs show data up to 12/28/10 at 1641 GMT. Stopped program and shut down spectrometer. Rebooted computer. Program and spectrometer appears to have started correctly. Notified mentor. RTS at 2150 GMT.